# Pseudo Transitions in the Finite-Size Blume-Capel Model

### v2

@@ Abstract
This article investigates the pseudo transitions of the Blume-Capel model on two-dimensional finite-size lattices. By employing the Wang-Landau sampling method and microcanonical inflection point analysis, we identified the positions of phase transitions as well as higher-order phase transitions. Through Metropolis sampling and canonical ensemble analysis, we obtained the corresponding geometric characteristics of the system at these transition points. The results indicate the presence of a third-order independent phase transition in the system. However, when the crystal field parameter $D$ exceeds 1.965, crossing the tricritical point, no third-order dependent phase transition is observed. Furthermore, the positions of the third-order phase transition obtained from both microcanonical and canonical analyses are consistent and mutually corroborative. We speculate that third-order dependent transitions may only occur in second-order phase transitions and not in first-order transitions.
@@
[label1]organization=College of Sciences, addressline=Xi’an University of Science and Technology, city=Xi’an, postcode=710054, country=China
@@
[label2]organization=School of Systems Science, addressline=Beijing Normal University, city=Beijing, postcode=100875, country=China
@@
[label3]organization=International Academic Center of Complex Systems, addressline=Beijing Normal University, city=ZhuHai, postcode=519087, country=China
@@
[label4]organization=Department of Systems Science, Faculty of Arts and Sciences, addressline=Beijing Normal University, city=ZhuHai, postcode=519087, country=China
@@
[label5]organization=2020 X-Lab, Shanghai Institute of Microsystem and Information Technology, addressline=Chinese Academy of Sciences, city=ShangHai, country=China
@@ 1 Introduction
Phase transitions are a common physical phenomenon in nature and constitute a central topic in classical statistical physics and thermodynamics. Phase transitions represent the most dramatic collective phenomena of many-body systems [1] and can alternatively be understood as sudden changes of the topology of the N-body configuration space [2, 3]. Such abrupt transitions can lead to catastrophic consequences, such as irreversible climate change [4, 5] or ecological collapse [6, 7]. Consequently, accurately identifying critical points and developing early warning systems have become major areas of research focus. Researchers have found that systems exhibit characteristic behavior in temporal [8, 9, 10] and spatial [11, 12, 13] dimensions as they approach critical transitions. Some researchers have also investigated early warning signals of phase transitions using methods that include thermodynamic quantities [14] and machine learning techniques [15]. Our objective is to identify characteristic indicators of early warning signals for phase transitions in spin systems, which can provide a foundation for critical early warnings and potentially be applied to real-world systems.
In recent developments, microcanonical analysis has increasingly been utilized for identifying phase transitions [1, 16]. This method was further generalized by Qi and Bachmann [17] to identify higher-order transitions, where independent and dependent transitions were distinguished. Independent transitions are similar to traditional phase transitions, occurring independently of other processes within the system. Dependent transitions rely on the occurrence of lower-order transitions. Independent and dependent transitions occur on either side of the phase transition point, indicating potential for early warning signals of phase transitions. By applying this method, researchers have demonstrated their effectiveness in the Ising [18], Potts [19], and Baxter-Wu models [20], including the detection of signals indicative of third-order transitions. The use of these analytical tools has revealed early signs of phase transitions, offering a strong basis for predicting critical shifts in spin systems. A recent study on ice premelting by Tian et al.[21] has attracted significant attention. Ice premelting occurs when water molecules migrate from beneath the ice surface to the surface, disrupting the initially ordered structure. This disturbance is analogous to the way isolated spins perturb the ordered state in spin systems. The discovery of this phenomenon demonstrated the existence of third-order transitions in real-world systems, thereby further validating the theoretical framework. This discovery not only deepened the understanding of third-order transitions, but also introduced new perspectives for future research. Building on these foundations, we aim to explore phase transition behavior through the Blume-Capel model.
@@
Exploring the changes in higher-order phase transitions on both sides of a system’s transition from second-order to first-order is an intriguing problem. Specifically, understanding how the system’s geometric characteristics evolve during higher-order phase transitions will be invaluable for developing effective early warning strategies for phase transitions in the future. Building on this, the density of states (DOS) of the Blume-Capel model was obtained using the efficient Wang-Landau algorithm [47, 48, 49, 50, 51, 52, 53], while the geometric properties of the Blume-Capel model were examined through Metropolis sampling. Building on these data, we analyzed phase transition behavior using both microcanonical and canonical methods and identified the locations of higher-order transitions. In the second section, the model and analytical methods are introduced. The third section presents and discusses the results, and the final section offers a summary and future perspectives on our work.
@@ 2.2 Wang-Landau Sampling
The density of states $g(E)$ is estimated through a random walk in energy space, with the histogram flattening as a key criterion, using the Wang-Landau (WL) sampling method[48]. As the exact values of the density of states are unknown at the beginning of the simulation, all entries are initially set to $g(E)=1$ and subsequently refined according to a specified probability distribution.
@@
where $E$ represents the energy level of the accepted state and $f$ is the modification factor, initially set to $f=f_{0}=e=2.71828$. Simultaneously, the energy histogram $H(E)$ is incremented by 1. The random walks continue, flipping spin values in the lattice configuration to generate new energy values for each iteration, until the energy histogram becomes ’flat’. A flat histogram is defined as the minimum value of any energy level’s histogram being at least 80% of the average value. The modification factor is then reduced to $f_{i+1}=\sqrt{f_{i}}$, and the histogram is reset to zero for the next iteration process. The simulation concludes when $f$ falls below the threshold of $f_{\text{final}}=1+10^{-8}$.
@@ 2.4 Geometrical analysis methods
To attain a more comprehensive understanding of this model, we conducted a further analysis from a geometric perspective. The spin configuration of the system was generated using Metropolis sampling [55]. The geometric order parameter, defined to characterize the system’s phase transition, was subsequently extracted from the configuration data. The first order parameter we used is the isolated spin.Our definition of isolated spins differs slightly from that in the work of Sitarachu and Bachmann [18]. In this study, an isolated spin is defined by two conditions: first, the spin’s state must differ from all its nearest neighbors; second, these neighboring spins must belong to the same cluster. Only when both conditions are met can the spin be classified as an isolated spin. The role of isolated spins is to disrupt both ordered and disordered states; thus, they may also be referred to as ’disruptors’. In the vicinity of the third-order independent transition, the number of isolated spins tends to increase. In this paper, we denote $I_{SO}$ to represent isolated spins. The second quantity is the average perimeter of clusters. Initially, we attempted to study the average cluster size, but we were unable to extract information related to the third-order dependent transition. We hypothesized that analyzing information from different dimensions within the same system would provide distinct insights into the phase transition. However, given that our model is constrained to a two-dimensional plane, we were unable to increase its dimensionality. As a result, we chose to apply dimensional reduction, focusing on the average perimeter of clusters. Using this approach, we successfully extracted signals associated with the third-order dependent transition.
@@ 3.1 Traditional transitions
where $\langle E^{n}\rangle=\sum_{E}E^{n}g(E)e^{-E/k_{B}T}$ and $N=L\times L$. We divide the specific heat by $N$ to enable the plotting of specific heat graphs for different sizes on the same chart, thereby facilitating our observation. In Fig. 2, we present a plot determining the phase transition points using specific heat for $D=1.0$ and $D=1.965$ in the canonical ensemble. We clearly observe that the specific heat exhibits a peak, with the temperature corresponding to this peak indicating the location of the phase transition point. To verify the accuracy of our data, we conducted a finite-size analysis. However, during the simulation, we approximated the energy level spacing to be 1, which resulted in the loss of some detailed information. Consequently, we corrected the specific heat in the first-order phase transition region using the following correction formula [56]:
@@
Figure 2: (Specific heat graph obtained under the canonical ensemble). Here, the specific heat calculated for different $D$ values is presented. Figure (a) shows the specific heat analysis for $D=1.0$, while Figure (b) illustrates the specific heat analysis for $D=1.965$. Different colors and markers are used to distinguish various system sizes.
Here, $\omega$ denotes the primary correction exponent, while $a$ and $b$ are non-universal constants. This correction enhanced the accuracy of the specific heat in the first-order phase transition, thereby increasing the reliability of the results. In the second-order phase transition region, the classical 2D Blume-Capel model belongs to the same universality class as the 2D Ising model. Consequently, we conducted a single-logarithmic analysis, as shown in Figure 3(a), which indicates that the system exhibits logarithmic divergence. However, in the first-order phase transition region, the critical behavior deviates significantly, and the system no longer follows the universality class of the 2D Ising model. According to Ref.[28], the expected critical exponent at the tricritical point is $\alpha/\nu=8/5$. The figure in Ref.[26] shows that the critical exponent near the tricritical point is approximately $\alpha/\nu\approx 1.608$. These reference values provide important theoretical support for our analysis. In this study, when $D=1.965$, the critical exponent obtained from Fig. 3(b) was $\alpha/\nu=1.57(1)$, which is in close agreement with the reference values, further validating the accuracy of our data and the reliability of our analytical methods.
Figure 3: (Finite size analysis of the Blume-Capel Model). In Fig. (a), a log-log analysis was conducted in the first-order phase transition region, with the extracted critical exponent added to the legend. In Fig. (b), a semi-logarithmic analysis was performed in the second-order phase transition region, and the extracted slope was similarly added to the legend.
@@
Figure 4: (For the Blume-Capel model with $L=50$, the canonical distribution $P(E,T)\sim g(E)e^{-E/k_{B}T}$ is examined at the transition temperature $T$). (a) For $D=1.0$, (b) For $D=1.965$. The red points represent data corresponding to the phase transition temperature, the black points represent data slightly below it, and the blue points represent data slightly above it.
@@
3.2 Third order transitions
@@
Figure 5: (Microcanonical inflection point analysis of the derivative data for D=1.0). The detailed results for the Blume-Capel model within the microcanonical ensemble at $D=1.0$ are presented. Figure (a) shows the variation of the first derivative of entropy, $\beta(E)$, while Figure (b) depicts the trend of the second derivative, $\gamma(E)$. Figure (c) highlights the evolution of the third derivative, $\delta(E)$, and Figure (d) reveals the changes in the fourth derivative, $\epsilon(E)$. To provide a clearer presentation of the results for different system sizes, different colors and shapes were carefully used to make each data point more distinct.
@@
In Fig. 5, we present the analysis for $D=1.0$. Figures 5(a)-(d) respectively display the first, second, third, and fourth derivatives of the density of states. In Fig. 5(a), a least-sensitive inflection point is observed, indicating the phase transition point under these conditions. In Fig. 5(b), a significant negative peak is observed in the region, indicating that the phase transition is of second order. In Fig. 5(c), we can observe a positive minimum and a negative maximum over the entire interval, representing signals for independent and dependent third-order transitions, respectively. Based on the analysis of Figure 5, we conclude that the phase transition is second order, and the signals for third-order transitions, along with their corresponding temperature and energy information, have been identified.
@@
Figure 6: (Microcanonical inflection point analysis of the derivative data for D=1.965). The detailed results for the Blume-Capel model within the microcanonical ensemble at $D=1.965$ are presented. Figure (a) shows the variation of the first derivative of entropy, $\beta(E)$, while Figure (b) depicts the trend of the second derivative, $\gamma(E)$. Figure (c) highlights the evolution of the third derivative, $\delta(E)$, and Figure (d) reveals the changes in the fourth derivative, $\epsilon(E)$. To provide a clearer presentation of the results for different system sizes, different colors and shapes were carefully used to make each data point more distinct.
@@
In Fig. 6, we conduct an in-depth analysis of the region near the tricritical point. A reference line at $y=0$ is added to Figs. 6 (b) and (c) to clarify the trends. In Fig. 6 (b), we observe a positive peak, indicating that the system is in the first-order transition region. It is observed that the peak value of $\gamma$ is relatively small, indicating that the system has just transitioned from the second-order phase transition region to the first-order phase transition region. In Fig. 6 (c), we observe two distinct positive minima on the left side of the transition point. However, as the system size increases, the minimum near the ground-state energy flattens and shows signs of disappearing. This suggests that the appearance of this point is likely due to noise in the simulation and boundary effects. The fluctuations observed in Fig. 6 (d) further support this hypothesis, providing strong evidences for our interpretation.
@@
Figure 7: (Microcanonical inflection point analysis of the derivative data for D=1.97). The detailed results for the Blume-Capel model within the microcanonical ensemble at $D=1.97$ are presented. Figure (a) shows the variation of the first derivative of entropy, $\beta(E)$, while Figure (b) depicts the trend of the second derivative, $\gamma(E)$. Figure (c) highlights the evolution of the third derivative, $\delta(E)$, and Figure (d) reveals the changes in the fourth derivative, $\epsilon(E)$. To provide a clearer presentation of the results for different system sizes, different colors and shapes were carefully used to make each data point more distinct.
@@
In Fig. 7, there is a certain distance from the tricritical point. As shown in Fig. 7(b), the height of the positive peak at this point is significantly higher than the peak in Fig. 6, indicating that the phase transition characteristics of the system are more pronounced. Similar to Figs. 6(c) and 7(c), we observe only a positive minimum and do not detect any signal of a negative maximum. This suggests that the phase transition trend in the system is becoming more uniform, and no signal for the third-order dependent transition exists in the first-order transition region. By comparing Figs. 5(b), 6(b), and 7(b), we observe that as the value of $D$ and system size $L$ increase, the peak of $\gamma(E)$ increases progressively but eventually converges to a fixed value. In fact, $\gamma(E)$ is related to a thermodynamic property associated with the specific heat phase transition. It is observed that for $D=1.0$, the peak of $\gamma(E)$ changes very little with system variation, whereas for $D=1.97$, the peak changes significantly but eventually approaches a fixed value. This occurs because a first-order phase transition is accompanied by latent heat, and as the system size increases, the latent heat gradually approaches the thermodynamic limit, eventually converges to a stable value. We summarize the analysis results for all parameters in Table 3. In the table, $1/\beta_{d}$ represents the temperature of the third-order dependent transition, while $1/\beta_{i}$ denotes the temperature of the third-order independent transition. ’NF’ indicates cases where no signal was found. The corresponding energy levels are also provided in the table to clearly illustrate the correlation between each signal and the system’s energy states.
@@
In Figs. 8 (a) and (b), we observe that the behavior of isolated spins is continuous within the second-order phase transition region. Thus, we define the maximum number of isolated spins as representing the position of the third-order independent transition, marked by black lines in the figs. 8 (a) and (b). By applying this approach, we determined the positions of the third-order independent transitions for $D=1.0$ and $D=1.5$ to be $T=1.207$ and $T=0.968$, respectively. Using microcanonical inflection point analysis, the corresponding values were found to be $T=1.3641$ and $T=1.1139$. The temperature difference of about 0.2 between the third-order independent transitions obtained by the two methods can be attributed to differences in the application of canonical and microcanonical methods, along with the discrete nature of energy levels (with an interval of 1) in the microcanonical analysis. These two factors together contribute to the observed systematic error.
@@
In Fig. 8 (c) and (d), we present the behavior of isolated spins at $D=1.965$ and $D=1.97$, where distinct discontinuous jumps in the number of isolated spins are observed. In Fig. 8 (c) and (d) insets provide magnified views of the regions where these discontinuities occur, with black and blue vertical lines marking the positions where the isolated spins begin to exhibit jumps and where the rate of change reaches its maximum. This jump phenomenon occurs because, as the crystal field parameter $D$ reaches a certain value, bringing the system closer to the phase transition, the fraction of spins in the 0 state rapidly increases, causing spins and clusters in the 0 state to dominate the system. In this scenario, the energy required to break clusters in the 0 state is lower than that required to break clusters in the -1 and 1 states. This is because the energy required for a spin to flip from the 0 state to either -1 or 1 is smaller than the energy required for a spin to flip from +1 to -1 (or from -1 to +1). As a result, even under minimal temperature variations, the number of isolated spins exhibits a suddenly increase, and at the point where the increase rate peaks, the ordered state of the system is disrupted, leading to a phase transition. Therefore, we define the position at which the isolated spins undergo discontinuous jumps as the signal of the third-order independent transition, and the position where the change in isolated spins is most pronounced as the phase transition point. In the first-order phase transition region, the temperature of the third-order independent transition signal is found to be very close to that of the phase transition point. However, as shown in the insets of Figs. 8 (c) and (d), the isolated spins exhibit two distinct states with a considerable difference between them. Near the first-order phase transition point, there exists a two-phase coexistence region with minimal temperature variation, yet both phases coexist within the system. This phenomenon is analogous to the coexistence of ice and water observed at the macroscopic scale, where two phases are present simultaneously. In microcanonical analysis, we came to the same conclusion, as shown in Table 3.

### v3

@@ Abstract
This article investigates the pseudo transitions of the Blume-Capel model on two-dimensional finite-size lattices. By employing the Wang-Landau sampling method and microcanonical inflection point analysis, we identified the positions of phase transitions as well as higher-order phase transitions. Through Metropolis sampling and canonical ensemble analysis, we determined the geometric characteristics of the system at these transition points. When the crystal field parameter $D$ exceeds 1.965, crossing the tricritical point, no third-order dependent phase transition is observed. However, a fourth-order independent transition was identified in the high-temperature region, and through Metropolis sampling analysis, a phase transition from the ordered paramagnetic phase to the disordered paramagnetic phase was confirmed, enhancing the phase diagram. Furthermore, the positions of the third-order phase transition obtained from both microcanonical and canonical analyses are consistent and mutually corroborative. We speculate that third-order dependent transitions vanish in the presence of strong first-order phase transitions.
@@
[label1]organization=College of Sciences, Xi’an University of Science and Technology, city=Xi’an, postcode=710054, country=China
@@
[label2]organization=National Innovation Institute of Defense Technology, city=Beijing, postcode=100071, country=China
@@
[label3]organization=School of Systems Science,Beijing Normal University, city=Beijing, postcode=100875, country=China
@@
[label4]organization=International Academic Center of Complex Systems, Beijing Normal University, city=Zhuhai, postcode=519087, country=China
@@
[label5]organization=Department of Systems Science, Faculty of Arts and Sciences, Beijing Normal University, city=Zhuhai, postcode=519087, country=China
[…]
[label6]organization=2020 X-Lab, Shanghai Institute of Microsystem and Information Technology, Chinese Academy of Sciences, city=Shanghai, postcode=200050, country=China
@@ 1 Introduction
Phase transitions are ubiquitous in nature and represent a fundamental topic in classical statistical physics and thermodynamics. They are among the most striking collective phenomena in many-body systems [1] and can also be interpreted as abrupt changes in the topology of the N-body configuration space [2, 3]. These sudden transitions can have catastrophic consequences, such as irreversible climate change [4, 5] or ecological collapse [6, 7]. Consequently, the precise identification of critical points and the development of early warning systems have become key research priorities. Studies have shown that systems exhibit distinct behaviors in both temporal [8, 9, 10] and spatial dimensions [11, 12, 13] as they approach critical transitions. Furthermore, early warning signals have been explored through approaches such as thermodynamic quantities [14] and machine learning-based methods [15]. In this study, we aim to identify characteristic indicators of early warning signals for phase transitions in spin systems, providing a foundation for critical early-warning methodologies and potential real-world applications.
In recent developments, microcanonical analysis has increasingly been utilized for identifying phase transitions [1, 16]. This method was further generalized by Qi and Bachmann [17] to identify higher-order transitions, where independent and dependent transitions were distinguished. Independent transitions are similar to traditional phase transitions, occurring independently of other processes within the system. Dependent transitions rely on the occurrence of lower-order transitions. Independent and dependent transitions occur on either side of the phase transition point, indicating potential for early warning signals of phase transitions. By applying this method, researchers have demonstrated their effectiveness in the Ising [18], Potts [19], and Baxter-Wu models [20], including the detection of signals indicative of third-order transitions. The use of these analytical tools has revealed early signs of phase transitions, offering a strong basis for predicting critical shifts in spin systems. A recent study on ice premelting by Tian et al [21] has attracted significant attention. Ice premelting occurs when water molecules migrate from beneath the ice surface to the surface, disrupting the initially ordered structure. This disturbance is analogous to the way isolated spins perturb the ordered state in spin systems. The discovery of this phenomenon demonstrated the existence of third-order transitions in real-world systems, thereby further validating the theoretical framework. This discovery not only deepened the understanding of third-order transitions, but also introduced new perspectives for future research. Building on these foundations, we aim to explore phase transition behavior through the Blume-Capel model.
@@
Exploring the changes in higher-order phase transitions across a system’s transition from second-order to first-order phases is an intriguing problem. Specifically, understanding how the system’s geometric characteristics evolve during higher-order phase transitions is crucial for devising effective early warning strategies in complex systems, including artificial swarms [47] and active bacterial colonies [48, 49]. The density of states (DOS) of the Blume-Capel model was obtained using the efficient Wang-Landau algorithm [50, 51, 52, 53, 54, 55, 56], while the geometric properties of the Blume-Capel model were examined through Metropolis sampling. Using these data, we examined phase transition behavior through both microcanonical and canonical methods, identifying the locations of higher-order transitions. In the second section, the model and analytical methods are introduced. The third section presents and discusses the results, and the final section provides a summary and outlines future research directions.
@@ 2.2 Wang-Landau Sampling
The Wang-Landau (WL) algorithm [51] is a highly efficient and broadly applicable method for directly estimating the density of states, $g(E)$, for a system. Utilizing $g(E)$ obtained through this method enables the direct investigation of higher-order phase transitions while avoiding additional processing, thereby minimizing errors and enhancing result accuracy. The density of states $g(E)$ is estimated through a random walk in energy space, with the histogram flattening as a key criterion, using the Wang-Landau (WL) sampling method. As the exact values of the density of states are unknown at the beginning of the simulation, all entries are initially set to $g(E)=1$ and subsequently refined according to a specified probability distribution.
@@
where $E$ represents the energy level of the accepted state and $f$ is the modification factor, initially set to $f=f_{0}=e=2.71828$. Simultaneously, the energy histogram $H(E)$ is incremented by 1. The random walks continue, flipping spin values in the lattice configuration to generate new energy values for each iteration, until the energy histogram becomes “flat”. A flat histogram is defined as the minimum value of any energy level’s histogram being at least 80% of the average value. The modification factor is then reduced to $f_{i+1}=\sqrt{f_{i}}$, and the histogram is reset to zero for the next iteration process. The simulation concludes when $f$ falls below the threshold of $f_{\text{final}}=1+10^{-8}$.
@@ 2.4 Geometrical analysis methods
To attain a more comprehensive understanding of this model, we conducted a further analysis from a geometric perspective. The spin configuration of the system was generated using Metropolis sampling [58]. The geometric order parameter, defined to characterize the system’s phase transition, was subsequently extracted from the configuration data. The first order parameter we used is the isolated spin. Our definition of isolated spins differs slightly from that in the work of Sitarachu and Bachmann [18]. In this study, an isolated spin is defined by two conditions: first, the spin’s state must differ from all its nearest neighbors; second, these neighboring spins must belong to the same cluster. Only when both conditions are met can the spin be classified as an isolated spin. The role of isolated spins is to disrupt both ordered and disordered states; thus, they may also be referred to as “disruptors”. In the vicinity of the third-order independent transition, the number of isolated spins tends to increase. In this paper, we denote $I_{SO}$ to represent isolated spins. The second quantity is the average perimeter of clusters. Initially, we attempted to study the average cluster size, but we were unable to extract information related to the third-order dependent transition. We hypothesized that analyzing information from different dimensions within the same system would provide distinct insights into the phase transition. However, given that our model is constrained to a two-dimensional plane, we were unable to increase its dimensionality. As a result, we chose to apply dimensional reduction, focusing on the average perimeter of clusters. Using this approach, we successfully extracted signals associated with the third-order dependent transition.
@@ 3.1 Traditional transitions
where $\langle E^{n}\rangle=\sum_{E}E^{n}g(E)e^{-E/k_{B}T}$ and $N=L\times L$. We divide the specific heat by $N$ to enable the plotting of specific heat graphs for different sizes on the same chart, thereby facilitating our observation. In Fig. 2, we present a plot determining the phase transition points using specific heat for $D=1.0$ and $D=1.965$ in the canonical ensemble. We clearly observe that the specific heat exhibits a peak, with the temperature corresponding to this peak indicating the location of the phase transition point. To verify the accuracy of our data, we conducted a finite-size analysis. However, during the simulation, we approximated the energy level spacing to be 1, which resulted in the loss of some detailed information. Consequently, the specific heat was adjusted using the following formula [59] when the primary phase transition of the system is of the first order:
@@
Figure 2: (Specific heat graph obtained under the canonical ensemble). The specific heat calculated at various $D$ values is presented here, with the peak of the specific heat serving as an indicator of the phase transition point. Figure (a) shows the specific heat analysis for $D=1.0$, while Figure (b) illustrates the specific heat analysis for $D=1.965$. Different colors and markers are used to distinguish various system sizes.
Here, $\omega$ denotes the primary correction exponent, while $a$ and $b$ are non-universal constants. This correction enhanced the accuracy of the specific heat in the first-order phase transition, thereby increasing the reliability of the results. When the primary phase transition of the system is of the second order, the classical 2D Blume-Capel model is categorized within the same universality class as the 2D Ising model. Consequently, we conducted a single-logarithmic analysis, as shown in Figure 3(a), which indicates that the system exhibits logarithmic divergence. However, during first-order phase transitions, the critical behavior markedly deviates, and the system ceases to adhere to the universality class of the 2D Ising model. According to Ref. [28], the expected critical exponent at the tricritical point is $\alpha/\nu=1.60$. The figure in Ref. [26] shows that the critical exponent near the tricritical point is approximately $\alpha/\nu\approx 1.608$. These reference values provide important theoretical support for our analysis. In this study, when $D=1.965$, the critical exponent obtained from Fig. 3(b) was $\alpha/\nu=1.57(1)$, which is in close agreement with the reference values, further validating the accuracy of our data and the reliability of our analytical methods.
Figure 3: (Finite size analysis of the Blume-Capel Model). Figure (a) displays a semi-logarithmic analysis conducted when the primary phase transition of the system is of the second order, revealing that its critical behavior follows logarithmic divergence, where $k$ represents the extracted slope. Figure (b) illustrates a log-log analysis performed when the primary phase transition is of the first order, incorporating the extracted critical exponent into the legend, where the tricritical point is characterized by a critical exponent of $\alpha/\nu=1.57$.
@@
Figure 4: (For the Blume-Capel model with $L=50$, the canonical distribution $P(E,T)\sim g(E)e^{-E/k_{B}T}$ is examined at the transition temperature $T$). (a) For $D=1.0$, the distribution of $P$ displays a single-peak form, suggesting that the primary phase transition is of the second order. (b) For $D=1.965$, the distribution of $P$ shows a double-peak form, implying that the primary phase transition is of the first order. The red points represent data corresponding to the phase transition temperature, the black points represent data slightly below it, and the blue points represent data slightly above it.
@@
3.2 Higher order transitions
@@
Figure 5: (Microcanonical inflection point analysis of the derivative data for D=1.0). The detailed results for the Blume-Capel model within the microcanonical ensemble at $D=1.0$ are presented. Figure (a) shows the variation of the first derivative of entropy, $\beta(E)$, while Figure (b) depicts the trend of the second derivative, $\gamma(E)$. Figure (c) highlights the evolution of the third derivative, $\delta(E)$, with a positive minimum and a negative maximum identified as indicators of third-order independent and dependent transitions, and Figure (d) reveals the changes in the fourth derivative, $\epsilon(E)$.
@@
In Fig. 5, we present the analysis for $D=1.0$. Figures 5(a)-(d) respectively display the first, second, third, and fourth derivatives of the density of states. In Fig. 5(a), a least-sensitive inflection point is observed, indicating the phase transition point under these conditions. In Fig. 5(b), a significant negative peak is observed in the region, indicating that the phase transition is of second order. In Fig. 5(c), we can observe a positive minimum and a negative maximum over the entire interval, representing signals for independent and dependent third-order transitions, respectively. Based on the analysis of Figure 5, we conclude that when $D=1.0$, the primary phase transition in the system is of second order, and the signals for third-order transitions, along with their corresponding temperature and energy information, are also identified.
@@
Figure 6: (Microcanonical inflection point analysis of the derivative data for D=1.965). The detailed results for the Blume-Capel model within the microcanonical ensemble at $D=1.965$ are presented. Figure (a) shows the variation of the first derivative of entropy, $\beta(E)$, while Figure (b) depicts the trend of the second derivative, $\gamma(E)$. Figure (c) highlights the evolution of the third derivative, $\delta(E)$, with a positive minimum at higher energy levels identified as an indicator of a third-order independent transition, and Figure (d) reveals the changes in the fourth derivative, $\epsilon(E)$.
@@
In Fig. 6, we conduct an in-depth analysis of the region near the tricritical point. A reference line at $y=0$ is added to Figs. 6 (b) and (c) to clarify the trends. In Fig. 6(b), we observe a positive peak, indicating that the primary phase transition of the system is first-order. The observed peak value of $\gamma$ is relatively small, suggesting that the system’s primary phase transition recently shifted from second-order to first-order. In Fig. 6 (c), we observe two distinct positive minima on the left side of the transition point. However, as the system size increases, the minimum near the ground-state energy flattens and shows signs of disappearing. This suggests that the appearance of this point is likely due to noise in the simulation and boundary effects. The fluctuations observed in Fig. 6 (d) further support this hypothesis, providing strong evidences for our interpretation. Therefore, the minima observed at higher energy levels are identified as indicators of third-order independent transitions.
@@
Figure 7: (Microcanonical inflection point analysis of the derivative data for D=1.97). The detailed results for the Blume-Capel model within the microcanonical ensemble at $D=1.97$ are presented. Figure (a) shows the variation of the first derivative of entropy, $\beta(E)$, with an inset depicting its extended part. Figure (b) depicts the trend of the second derivative, $\gamma(E)$. Figure (c) illustrates the evolution of the third derivative, $\delta(E)$, with the positive minimum identified as an indicator of a third-order independent transition, and Figure (d) showcases the changes in the fourth derivative, $\epsilon(E)$, with an inset showing its extended part and an orange vertical line indicating the identified position of the fourth-order independent transition.
@@
In Fig. 7, there is a certain distance from the tricritical point. As shown in Fig. 7(b), the height of the positive peak at this point is significantly higher than the peak in Fig. 6, indicating that the phase transition characteristics of the system are more pronounced. Similar to Figs. 6(c) and 7(c), we observe only a positive minimum and do not detect any signal of a negative maximum. This suggests that when the primary phase transition of the system is of the first order, there is no signal for a third-order dependent transition. However, our exploration of the high-temperature region has revealed a signal for a fourth-order independent transition, indicated by an orange vertical line in the inset of Figure 7(d), complemented by a corresponding inset in Figure 7(a) to facilitate temperature determination. By comparing Figs. 5(b), 6(b), and 7(b), we observe that as the value of $D$ and system size $L$ increase, the peak of $\gamma(E)$ increases progressively but eventually converges to a fixed value. In fact, $\gamma(E)$ is related to a thermodynamic property associated with the specific heat phase transition. It is observed that for $D=1.0$, the peak of $\gamma(E)$ changes very little with system variation, whereas for $D=1.97$, the peak changes significantly but eventually approaches a fixed value. This occurs because a first-order phase transition is accompanied by latent heat, and as the system size increases, the latent heat gradually approaches the thermodynamic limit, eventually converges to a stable value. We summarize the analysis results for all parameters in Table 3. In the table, $1/\beta_{d}$ represents the temperature of the third-order dependent transition, while $1/\beta_{i}$ denotes the temperature of the third-order independent transition. “NF” indicates cases where no signal was found. The corresponding energy levels are also provided in the table to clearly illustrate the correlation between each signal and the system’s energy states.
@@
In Figs.8(a) and (b), we observe that isolated spins exhibit continuous changes and peak as the system undergoes a primary phase transition of the second order. This phenomenon occurs as the temperature increases, leading to a gradual rise in the number of isolated spins, which consequently causes the system to fragment. However, once the system reaches a certain level of fragmentation, these isolated spins begin to aggregate into smaller clusters, thereby reducing the number of isolated spins, ultimately leading to the emergence of this peak. Thus, we define the peak of isolated spins as representing the position of the third-order independent transition, marked by black lines in Figs. 8 (a) and (b). By applying this approach, we determined the positions of the third-order independent transitions for $D=1.0$ and $D=1.5$ to be $T=1.207$ and $T=0.968$, respectively. Using microcanonical inflection point analysis, the corresponding values were found to be $T=1.3641$ and $T=1.1139$. The temperature difference of about 0.2 between the third-order independent transitions obtained by the two methods can be attributed to differences in the application of canonical and microcanonical methods, along with the discrete nature of energy levels (with an interval of 1) in the microcanonical analysis. These two factors together contribute to the observed systematic error.
@@
In Figs. 8(c) and (d), we observe that isolated spins exhibit a discontinuous jump after reaching a certain proportion when the system undergoes a first-order primary phase transition. In the insets of Figs. 8 (c) and (d), magnified views show the regions where these discontinuities occur, with black and blue vertical lines indicating the positions where the isolated spins begin to exhibit jumps and where the rate of change reaches its maximum, respectively. We identify these positions as those of the third-order independent transition and the phase transition point, respectively. We define these as the positions of the third-order independent transition and the phase transition point, respectively. This jump phenomenon occurs because, as the crystal field parameter $D$ reaches a certain value, bringing the system closer to the phase transition, the fraction of spins in the 0 state rapidly increases, causing spins and clusters in the 0 state to dominate the system. In this scenario, the energy required to break clusters in the 0 state is lower than that required to break clusters in the -1 and 1 states. This is because the energy required for a spin to flip from the 0 state to either -1 or 1 is smaller than the energy required for a spin to flip from +1 to -1 (or from -1 to +1). As a result, even with minimal temperature variations, the number of isolated spins increases abruptly. At the point where the rate of increase reaches its peak, a phase transition occurs in the system. Therefore, we define the position at which the isolated spins undergo discontinuous jumps as the signal of the third-order independent transition, and the position where the change in isolated spins is most pronounced as the phase transition point. When the primary phase transition is of the first order, we find that the temperatures of the third-order independent transition and the phase transition point are very close. However, as shown in the insets of Figs. 8 (c) and (d), the isolated spins exhibit two distinct states with a considerable difference between them. Similar conclusions were reached in our microcanonical inflection point analysis.